# Evaluating AI systems under uncertain ground truth: a case study in dermatology

### v1

@@ Appendix C Case study details
Figure 12: We plot the number of annotators per case and the number of unique conditions annotated by these annotators for low- (left) and high-risk (right) cases. Risk categories are based on the risk-level of the top-1 IRN condition as in previous work (Roy et al., 2022). Surprisingly, high-risk cases were annotated by, on average, less annotators while clearly yielding more unique conditions.
@@
We use the de-identified dermatology dataset previously used in (Liu et al., 2020) which was collected from $17$ different sites across California and Hawaii. Each case in the dataset consists of up to $3$ RGB images, taken by medical assistants using consumer-grade digital cameras in a clinical environment. The images exhibit a large amount of variation in terms of affected anatomic location, background objects, resolution, perspective and lighting. Our dataset has patients with different skin types varying between Fitzpatrick scale of $1-6$. We resized each image to $448\times 448$ pixels for our training and evaluation purposes. The dataset consists of $16,225$ cases, where $12,335$ cases are used for model training, $1951$ cases for validation and $1939$ cases for test set. Note that in our dataset there are a few patients with multiple cases collected during repeated visits over the duration of the data collection process. For the data split, we ensured a patient-level split to avoid any bias. Furthermore, we ensured a similar distribution of skin conditions and risk categories. The latter corresponds to a classification of case outcomes for a given skin condition, i.e., each skin condition has been mapped to a risk category in low, medium and high. We operate on a label space of $419$ skin conditions similar to (Liu et al., 2020), following a long-tailed distribution. The dataset was labelled by US or Indian board certified dermatologists with $5-30$ years of experience. Annotation was conducted over a period of time in tandem to data collection between $2010-2018$. As a result, different cases were annotated by different dermatologists (see Figure 12 for statistics). Moreover, there are varying numbers of annotations per case but we made sure to have at least $3$ annotations per case.
C.1 Annotation details
@@
All dermatologists were actively seeing patients in clinics and passed a certification test to ensure they were comfortable grading cases in our labeling tool. To provide the differential diagnoses, each annotator selected one or multiple conditions from the standardized Systematized Nomenclature of Medicine-Clinical Terms (SNOMED-CT) [^5] http://snomed.org/ using a search-as-you-type interface, and assigned a confidence rating per condition. The conditions were later mapped to the 419 conditions as described previously (Liu et al., 2020). The confidence ratings were then used to rank conditions since confidence assignments are not comparable across dermatologists. As multiple conditions can obtain the same confidence ratings, this results in the partial rankings described in Section B. These rankings are the basis for our study and several prior works (Eng et al., 2019; Roy et al., 2021; Azizi et al., 2022). As the associated risk levels are of particular interest for analysis in prior work (Roy et al., 2021), it is important to highlight that different annotator annotations also correspond to different risk annotations. Note that risk is not specific to each case but rather based on a fixed condition to risk mapping.
C.2 Training details
@@
As we have multiple images per case, we are dealing with a multi-instance classification task with $419$ labels. As we are dealing with limited training data, we use the BiT-L model (Kolesnikov et al., 2020), pre-trained on JFT dataset (Sun et al., 2017) containing large-scale natural images from the internet. The encoder provides a feature representation for each image instance in a case. This output is then passed through an instance level average pooling layer, which generates a common feature-map for all the instances for a given case. This is passed to a classification head through a drop-out layer. The classification head has an intermediate fully-connected hidden layer with $1024$-dimensions followed by a final classification with a softmax layer which provide probabilities for each of the $419$ classes. To keep model training process consistent to prior work (Liu et al., 2020; Roy et al., 2022), we use IRN as (soft) supervision in a cross-entropy loss. For data augmentation, we use horizontal and vertical flips, rotations, variations of brightness, contrast and saturation, and random Gaussian blurring. We use a Adam optimizer for training. Each model is trained for $10,000$ steps. Convergence of all the models were ensured on the validation set. We use a batch-size of $16$ cases for training. For each model, we select the final checkpoint based on the highest top-3 accuracy against the top-1 IRN ground truth label, consistent with previous work. We train a pool of 30 models by randomly varying the size of the ResNet architecture (depth $\times$ width: $101\times 3$, $101\times 1$, $50\times 3$, $152\times 2$, $152\times 4$), pre-logit dropout rate (keep probability: $0.5-0.7$) and learning rate ($10^{-6}-5\times 10^{-4}$). As model performance is not the focus of this work, we randomly selected four of these models for our experiments.
C.3 PL and PrIRN details
@@
As detailed in Section B.4, for PL, we repeat annotators in order to simulate higher reliability. Specifically, corresponding to the x-axis in Figure 6, we evaluate 1, 2, 3, 5, and 10 repetitions. For PrIRN, we evaluate temperatures $\gamma=10,20,30,50,100$.
Appendix D Additional results
@@
Figure 13: PL annotation certainty (blue) at medium reliability (see Figure 7) plotted with annotator agreement (red), see text for details. With a correlation coefficient of $0.85$, both measures are highly correlated, indicating that annotation certainty indeed captures the uncertainty induced by disagreement.
@@
Figure 14: Top-1 annotation certainty for PL (left) and PrIRN across reliabilities. As detailed in the text, for PL, reliability is approximated by repeating annotations; for PrIRN, we can directly increase the corresponding temperature parameter.
Reliabity in PL and PrIRN: Figure 14 plots top-1 annotation certainty for PL and PrIRN across reliabilities. As detailed in Section B, reliability can be understood as the number of annotators in both cases. For PL, we explicitly repeat annotators as integrating a temperature into the statistical aggregation model makes sampling from the posterior infeasible; for PrIRN, in contrast, Section B.3 allows the direct use of a temperature. As indicated by Figure 6 (left), average annotation certainty increases with reliability. This suggests that annotation certainty increases across examples. However, this is not entirely true; we clearly see a saturating effect for at least 250 cases where higher reliability does not seem to increase annotation certainty. This further supports our observation that the observed disagreement for many cases is actually due to inherent uncertainty, i.e., these cases are inherently ambiguous for annotators.
@@
Refer to caption A0: {Pyoderma gangrenosum, Venous stasis ulcer} A1: {Arterial ulcer, Calciphylaxis cutis} A2: {Cellulitis}
@@
Model A: $C_{\text{top-}3}=$Diabetic ulcer, Pyoderma Gangrenosum, Venous Stasis Ulcer} Model D: $C_{\text{top-}3}=$Arterial ulcer, Cellulitis, Pyoderma Gangrenosum}
@@
Model A PL PrIRN $\mathbb{E}_{p(\lambda|b,x)}\{\delta[Y_{\text{top-}1}(\lambda)\subset C_{\text{top-}k}(x)]\}$ $0.39$ $0.41$ Model D PL PrIRN $\mathbb{E}_{p(\lambda|b,x)}\{\delta[Y_{\text{top-}1}(\lambda)\subset C_{\text{top-}k}(x)]\}$ $0.58$ $0.61$
@@
Refer to caption A0: {Dissecting cellulitis of scalp, Tinea} {Seborrheic dermatitis} A1: {Acne keloidalis, Folliculitis decalvans} A2: {Folliculitis}
@@
Model A: $C_{\text{top-}3}=$Acne, Folliculitis, Folliculitis decalvans} Model B: $C_{\text{top-}3}=$Acne keloidalis, Folliculitis, Folliculitis decalvans}
@@
Model A PL PrIRN $\mathbb{E}_{p(\lambda|b,x)}\{\delta[Y_{\text{top-}1}(\lambda)\subset C_{\text{top-}k}(x)]\}$ $0.27$ $0.4$ Model B PL PrIRN $\mathbb{E}_{p(\lambda|b,x)}\{\delta[Y_{\text{top-}1}(\lambda)\subset C_{\text{top-}k}(x)]\}$ $0.42$ $0.62$
Figure 15: Two additional qualitative results corresponding to cases with very low annotation certainty, following the presentation of Figure 5.
@@
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" alt="[LOGO]">

### v2

@@ Appendix D Case study details
Fig. 13: We plot the number of annotators per case and the number of unique conditions annotated by these annotators for low- (left) and high-risk (right) cases. Risk categories are based on the risk-level of the top-1 IRN condition as in previous work [84]. Surprisingly, high-risk cases were annotated by, on average, less annotators while clearly yielding more unique conditions.
@@
We use the de-identified dermatology dataset previously used in [58] which was collected from $17$ different sites across California and Hawaii. Each case in the dataset consists of up to $3$ RGB images, taken by medical assistants using consumer-grade digital cameras in a clinical environment. The images exhibit a large amount of variation in terms of affected anatomic location, background objects, resolution, perspective and lighting. Our dataset has patients with different skin types varying between Fitzpatrick scale of $1-6$. We resized each image to $448\times 448$ pixels for our training and evaluation purposes. The dataset consists of $16,225$ cases, where $12,335$ cases are used for model training, $1951$ cases for validation and $1939$ cases for test set. Note that in our dataset there are a few patients with multiple cases collected during repeated visits over the duration of the data collection process. For the data split, we ensured a patient-level split to avoid any bias. Furthermore, we ensured a similar distribution of skin conditions and risk categories. The latter corresponds to a classification of case outcomes for a given skin condition, i.e., each skin condition has been mapped to a risk category in low, medium and high. We operate on a label space of $419$ skin conditions similar to [58], following a long-tailed distribution. The dataset was labelled by US or Indian board certified dermatologists with $5-30$ years of experience. Annotation was conducted over a period of time in tandem to data collection between $2010-2018$. As a result, different cases were annotated by different dermatologists (see Figure 13 for statistics). Moreover, there are varying numbers of annotations per case but we made sure to have at least $3$ annotations per case.
D.1 Annotation details
@@
All dermatologists were actively seeing patients in clinics and passed a certification test to ensure they were comfortable grading cases in our labeling tool. To provide the differential diagnoses, each annotator selected one or multiple conditions from the standardized Systematized Nomenclature of Medicine-Clinical Terms (SNOMED-CT) [^12] http://snomed.org/ using a search-as-you-type interface, and assigned a confidence rating per condition. The conditions were later mapped to the 419 conditions as described previously [58]. The confidence ratings were then used to rank conditions since confidence assignments are not comparable across dermatologists. As multiple conditions can obtain the same confidence ratings, this results in the partial rankings described in Section 2.3. These rankings are the basis for our study and several prior works [30, 84, 5]. As the associated risk levels are of particular interest for analysis in prior work [84], it is important to highlight that different annotator annotations also correspond to different risk annotations. Note that risk is not specific to each case but rather based on a fixed condition to risk mapping.
D.2 Training details
@@
As we have multiple images per case, we are dealing with a multi-instance classification task with $419$ labels. As we are dealing with limited training data, we use the BiT-L model [50], pre-trained on JFT dataset [101] containing large-scale natural images from the internet. The encoder provides a feature representation for each image instance in a case. This output is then passed through an instance level average pooling layer, which generates a common feature-map for all the instances for a given case. This is passed to a classification head through a drop-out layer. The classification head has an intermediate fully-connected hidden layer with $1024$-dimensions followed by a final classification with a softmax layer which provide probabilities for each of the $419$ classes. To keep model training process consistent to prior work [58, 84], we use IRN as (soft) supervision in a cross-entropy loss. For data augmentation, we use horizontal and vertical flips, rotations, variations of brightness, contrast and saturation, and random Gaussian blurring. We use a Adam optimizer for training. Each model is trained for $10,000$ steps. Convergence of all the models were ensured on the validation set. We use a batch-size of $16$ cases for training. For each model, we select the final checkpoint based on the highest top-3 accuracy against the top-1 IRN ground truth label, consistent with previous work. We train a pool of 30 models by randomly varying the size of the ResNet architecture (depth $\times$ width: $101\times 3$, $101\times 1$, $50\times 3$, $152\times 2$, $152\times 4$), pre-logit dropout rate (keep probability: $0.5-0.7$) and learning rate ($10^{-6}-5\times 10^{-4}$). As model performance is not the focus of this work, we randomly selected four of these models for our experiments.
D.3 PL and PrIRN details
@@
As detailed in Section 2.3.3, for PL, we repeat annotators in order to simulate higher reliability. Specifically, corresponding to the x-axis in Figure 6, we evaluate 1, 2, 3, 5, and 10 repetitions. For PrIRN, we evaluate temperatures $\gamma=10,20,30,50,100$.
Appendix E Additional results
@@
Fig. 14: PL annotation certainty (blue) at medium reliability (see Figure 7) plotted with annotator agreement (red), see text for details. With a correlation coefficient of $0.85$, both measures are highly correlated, indicating that annotation certainty indeed captures the uncertainty induced by disagreement.
@@
Fig. 15: Top-1 annotation certainty for PL (left) and PrIRN across reliabilities. As detailed in the text, for PL, reliability is approximated by repeating annotations; for PrIRN, we can directly increase the corresponding temperature parameter.
Reliabity in PL and PrIRN: Figure 15 plots top-1 annotation certainty for PL and PrIRN across reliabilities. As detailed in Section 2, reliability can be understood as the number of annotators in both cases. For PL, we explicitly repeat annotators as integrating a temperature into the statistical aggregation model makes sampling from the posterior infeasible; for PrIRN, in contrast, Section 2.3.2 allows the direct use of a temperature. As indicated by Figure 6 (left), average annotation certainty increases with reliability. This suggests that annotation certainty increases across examples. However, this is not entirely true; we clearly see a saturating effect for at least 250 cases where higher reliability does not seem to increase annotation certainty. This further supports our observation that the observed disagreement for many cases is actually due to inherent uncertainty, i.e., these cases are inherently ambiguous for annotators.
Fig. 16: Distributions of top-1 risk and condition certainty, complementary to Figure 8.
Top-1 risk and condition certanties: Figure 16 shows the distribution of top-1 risk and condition certainties, complementing Figure 8 in the main paper. Clearly, more cases exhibit high top-1 risk certainty compared to top-1 condition certainty because there are only three risk categories of 419 different conditions. However, the fraction of examples with certainty below 99% is with 38.2% still very high (compared to 59% for conditions).
A0: {Pyoderma gangrenosum, Venous stasis ulcer}
A1: {Arterial ulcer, Calciphylaxis cutis}
A2: {Cellulitis}
@@
Model A: $C_{\text{top-}3}=$
{Diabetic ulcer, Pyoderma Gangrenosum, Venous Stasis Ulcer}
Model D: $C_{\text{top-}3}=$
{Arterial ulcer, Cellulitis, Pyoderma Gangrenosum}
@@
Model A PL PrIRN
$\mathbb{E}_{p(\lambda|b,x)}\{\delta[Y_{\text{top-}1}(\lambda)\subset C_{\text{top-}k}(x)]\}$ $0.39$ $0.41$
Model D PL PrIRN
$\mathbb{E}_{p(\lambda|b,x)}\{\delta[Y_{\text{top-}1}(\lambda)\subset C_{\text{top-}k}(x)]\}$ $0.58$ $0.61$
@@
A0: {Dissecting cellulitis of scalp, Tinea} {Seborrheic dermatitis}
A1: {Acne keloidalis, Folliculitis decalvans}
A2: {Folliculitis}
@@
Model A: $C_{\text{top-}3}=$
{Acne, Folliculitis, Folliculitis decalvans}
Model B: $C_{\text{top-}3}=$
{Acne keloidalis, Folliculitis, Folliculitis decalvans}
@@
Model A PL PrIRN
$\mathbb{E}_{p(\lambda|b,x)}\{\delta[Y_{\text{top-}1}(\lambda)\subset C_{\text{top-}k}(x)]\}$ $0.27$ $0.4$
Model B PL PrIRN
$\mathbb{E}_{p(\lambda|b,x)}\{\delta[Y_{\text{top-}1}(\lambda)\subset C_{\text{top-}k}(x)]\}$ $0.42$ $0.62$
Fig. 17: Two additional qualitative results corresponding to cases with very low annotation certainty, following the presentation of Figure 5.
@@
Differences between PL and PrIRN: In Figure 18 we show further qualitative results highlighting some cases with differences between the PL and PrIRN plausibilities.
A0: {SK/ISK}
A1: {Actinic Keratosis}
A2: {SK/ISK}
A3: {Verruca vulgaris}
A4: {Lentigo}
A5: {Lentigo}
Annotations
Plausibilities
PL plausibilities                                                          PRIRN plausibilities
Refer to captionRefer to caption
Annot.
A0: {Cylindroma of skin, SK/ISK} {Cyst}
A2: {Nevus sebaceous}
A3: {Basal Cell Carcinoma}
A4: {Nevus sebaceous}
Plausibilities
PL plausibilities                                                          PRIRN plausibilities
Refer to captionRefer to caption
Annot.
A0: {Sweet syndrome}
A1: {Drug Rash, Eczema}
A2: {Actinic Keratosis} {Cutaneous lupus}
Plausibilities
PL plausibilities                                                          PRIRN plausibilities
Refer to captionRefer to caption
Annotations
A0: {Psoriasis} {Drug Rash}
A1: {Drug Rash, Eczema} {Psoriasis}
A2: {Psoriasis} {Cutaneous T Cell Lymphoma, Eczema}
A3: {Drug Rash, Eczema} {Cutaneous T Cell Lymphoma}
Plausibilities
PL plausibilities                                                          PRIRN plausibilities
Refer to captionRefer to caption
Fig. 18: Two additional qualitative results focusing on some differences between PL and PrIRN plausibilities.